OpenStudy (anonymous):

125^x+1 * 23^x+1=2500

OpenStudy (anonymous):

i have started it..may i know am i going right?? \[5^{3x+3} \times 23^{x+1} = 50^{2}\]

OpenStudy (wolf1728):

Going back to the beginning: That is 125^(x+1) * 23^(x+1) = 2500 correct?

OpenStudy (anonymous):

yaa correct @wolf1728

OpenStudy (wolf1728):

So, basically if x = .9825 then it equals 2500.9884533405

OpenStudy (anonymous):

how did you get that?? i cant understand??

OpenStudy (wolf1728):

Just doing a few calculations - trial and error. I always like to get a rough idea before doing any serious equation solving.

OpenStudy (wolf1728):

125^.9825 = 114.8719749473 23^.9825 = 21.7719635663 114.8719749473 * 21.7719635663 = 2500.9884533405

OpenStudy (vishweshshrimali5):

I think this is a logarithm question.

OpenStudy (vishweshshrimali5):

See: \[\large{125^{x+1} \times 23^{x+1} = 2500}\] \[\large{\implies 5^{3x+3} \times 23^{x+1} = 50^2}\] Taking log base 10 both sides, \[\large{\implies (3x+3)\log5 + (x+1)\log23 = 2\log50}\] \[\large{\implies (3x+3)\log5 + (x+1)\log23 = 2\log(5^2\times 2)}\] \[\large{\implies (3x+3)\log5 + (x+1)\log23 = 2(2\log 5 + \log2)}\] \[\large{\implies (3x+3)\log5 + (x+1)\log23 = 4\log5 + 2\log2}\] \[\large{\implies (3x-1)\log5 + (x+1)\log23 = 2\log2}\] \[\large{\implies 3x\log5 + x\log23 -\log5 + \log23 = 2\log2}\] \[\large{\implies x(3\log5+\log23) = 2\log2 + \log5 - \log23}\] \[\large{\implies x = \cfrac{2\log2+\log5 - \log23}{3\log5+\log23}}\] Now, using a calculator find out the value of x.

OpenStudy (vishweshshrimali5):

@dinisha , I hope this helps you.

OpenStudy (vishweshshrimali5):

Ask me if you don't understand any step.

OpenStudy (wolf1728):

I thought I was good at logarithms but that was one heck of a solution you did there!!

OpenStudy (vishweshshrimali5):

:)

OpenStudy (wolf1728):

I just ignored those "+1" terms in the exponents thinking they were totally inconsequential.

OpenStudy (wolf1728):

2*log(2) = 0.6020599913 log(5) = 0.6989700043 -log(23) = -1.361727836 Numerator SUM = -0.0606978404 3 * log(5) = 2.096910013 log(23) = 1.361727836 Denominator Sum = 3.458637849 Numerator / Denominator = -0.0175496375 (not quite .9825 is it???)

OpenStudy (vishweshshrimali5):

Perfectly correct calculation @wolf1728 :)

OpenStudy (vishweshshrimali5):

Even wolfram agrees with you. So, x = -0.0175... and that's correct :)

OpenStudy (wolf1728):

Oh geez - I just thought of the actual STUPID answer. After all those calculations and exponentiations, etc - I actually (with .9825) solved for X PLUS 1 . So the answer really is .9825 = X PLUS 1 or -.0175 How foolish of me to forget i was solving for X PLUS 1. Geez couldd they make this any more complicated ?????

OpenStudy (vishweshshrimali5):

:) Well I really have some sympathy for you That happens a lot..

OpenStudy (wolf1728):

Thanks - and thanks for the medal - even though you found the actual answer (and not the trial and error solution). Once again, good work vishweshshrimali5 !!!

OpenStudy (vishweshshrimali5):

Thanks and same to you @wolf1728. Trying is much more important than finding the correct answer. :) Good day

OpenStudy (wolf1728):

Guess so :-)

OpenStudy (anonymous):

but the answer provided are 5/4 or 1/4 or 3/4 or 2/3 or 1/3 .. even when i convert the answer into fraction its not similar to anything @vishweshshrimali5 and @wolf1728